# SI_appendix_phys

### v1

@@ B The method of coexisting phases and the binodal curve
C Linear stability of homogeneous solutions and the spinodal curve
The homogeneous solution $(\rho,\rho_{a},m)=(\phi,\phi_{a},0)$ is always a solution of Eqs. (S1)-(S3). We analyze the linear stability of this solution by taking $\rho=\phi+\delta\tilde{\rho}$ , $\rho_{a}=\phi_{a}+\delta\tilde{\rho}_{a}$ , $m=0+\delta\tilde{m}$. At linear order in $\delta$, we obtain
@@ C Linear stability of homogeneous solutions and the spinodal curve
This matrix is not Hermitian, so its eigenvalues are, in general, complex. However, if the spectrum is complex, then two of the eigenvalues form a complex conjugate pair and the other remains real; this is due to symmetry under $\lambda\mapsto\bar{\lambda}$ and $(A_{1},A_{2},A_{3})\mapsto(\bar{A}_{1},\bar{A}_{2},-\bar{A}_{3})$ in (S22).
As usual, the homogeneous state is stable if all eigenvalues $\lambda$ have negative real parts. (It is implicit that $\lambda$ depends on $q$; this condition must hold for all $q$.) The resulting instabilities can have several types; we classify them according to the scheme of Ref. [47]. Our system has two conserved densities $\rho,\rho_{a}$ and a non-conserved magnetization $m$. The behavior is controlled by the conserved densities, which restricts the behavior to four of the eight types considered in Ref. [47]. If the dominant eigenvalue of $W$ is real then they are called stationary, else they are called oscillatory; if the instability is initiated by modes with $q\to 0$ then it is called large scale, else it is called small scale. The resulting types are then conserved-Turing (stationary, small scale), Cahn–Hilliard (stationary, large scale), conserved-Hopf (oscillatory, large scale), or conserved-wave (oscillatory, small scale).
Fig. S1 illustrates the range of possible behavior, showing several instabilities that occur on increasing $\phi_{a}$ at fixed total density $\phi$. Row (a) shows the onset of a Cahn-Hilliard instability (stationary, large-scale), row (b) shows the onset of a conserved-Hopf instability (oscillatory, large-scale), and row (c) shows the onset of a conserved-wave instability (oscillatory, small-scale). In the left (right) column, the active volume fraction is sub(super)-critical $\phi_{a}<\phi_{a}^{*}$ $\max_{q}\text{Re}\lambda<0$ ($\phi_{a}>\phi_{a}^{*}$ $\max_{q}\text{Re}\lambda>0$). The central column displays the critical active volume fraction $\phi_{a}=\phi_{a}^{*}$, $\max_{q}\text{Re}\lambda/q^{2}=0$. We show below that these are the only possible scenarios because the onset of a stationary instability must occur on a large scale.
Figure S1: Eigenvalues of $W$ ((S27)) at the onset of instability. Active volume fraction $\phi_{a}$ increases from left to right. The critical volume fraction is displayed in the central column. (a) Cahn-Hiliard instability (stationary, large-scale). (b) Conserved-Hopf instability (oscillatory, large-scale). (c) Conserved-wave instability (oscillatory, small-scale).
@@ 1 Spinodal curves
At the boundary of linear stability, then $\text{Re}(\lambda)=0$ for at least one eigenvalue. There are two situations where this can happen. (i) $\text{det}(W)=0$, corresponding to a vanishing eigenvalue $\lambda=0$ (stationary instability). (ii) the characteristic polynomial is of the form $(\lambda^{2}+F(W))(\lambda-\text{Tr}(W))=0$, with $F(W)>0$ so that $\lambda=\pm i\sqrt{F(W)}$ is pure imaginary. This situation holds if and only if $F(W)\text{Tr}(W)=\text{det}(W)<0$ [we have always $\text{Tr}(W)<0$ so the final inequality ensures $F(W)>0$.].
In the stationary case one may solve $\det W=0$ for $\phi_{a}$ to obtain
@@
In the oscillatory case, one may similarly solve $\text{det}(W)=F(W)\text{Tr}(W)$ to obtain
$\phi_{a}^{{\rm osc}}(q)=\frac{2\phi\left(d_{s}(\phi)q^{2}+1\right)\left(2+2d_{s}(\phi)+d_{s}(\phi)^{2}\text{Pe}^{2}+(d_{s}(\phi)+1)^{2}q^{2}\right)}{\text{Pe}^{2}\left(-d_{s}(\phi)(1-\phi)\left(q^{2}(d_{s}^{\prime}(\phi)\phi+1)+2\right)-d_{s}^{\prime}(\phi)(1-\phi)\phi\left(2+q^{2}\right)+2d_{s}(\phi)^{3}q^{2}-d_{s}(\phi)^{2}\left(2+(1-\phi)q^{2}\right)\right)}.$ (S27)
@@
[For a finite system, the trial solution (S21) is restricted to $q=2n\pi/L$ with $n\in\mathbb{Z}$ and we should minimize over this discrete set, so $\phi_{a}^{*}$ depends in general on $L$. We consider the limit $L\to\infty$ here, so we take an infimum over $q>0$.] Note also: if the infimum in (S28) is achieved by $\phi_{a}^{{\rm osc}}(q)$ then it is certain that $\det(W)<0$ at this point, as required for an oscillatory instability: this holds because $\det(W)$ only changes sign at $\phi_{a}=\phi_{a}^{\rm st}(q)>\phi_{a}^{\rm osc}(q)$ and $\det(W)<0$ for $\phi_{a}\to 0$.
@@
We also observe from (S26) that $\inf_{q}\phi_{a}^{{\rm st}}(q)=\phi_{a}^{{\rm st}}(0)$, which means that if this instability is of stationary type then it is always large-scale, as already asserted above. On the other hand, the infimum of $\phi_{a}^{{\rm osc}}(q)$ may occur as $q\to 0$ (large-scale oscillatory instability) or at finite $q$ (small-scale oscillatory instability).
For numerical calculations, it is convenient to parameterize the dependence on $\phi_{a}$ in terms of the quantity $\nu$ defined in (5). Then $\nu=\frac{\phi}{1-\phi}$ corresponds to a system of purely passive particles and $\nu=0$ to purely active particles. Fig. S2 illustrates the spinodal in the $(\phi,\nu)$-plane. We consider a narrow range of Pe in which both oscillatory and stationary instabilities occur, and we plot separately the curves corresponding to $\phi_{a}=\inf_{q}\phi_{a}^{\rm osc}(q)$ and $\phi_{a}=\phi_{a}^{\rm st}(q)$. The boundary of linear instability is given by the smaller of these $\phi_{a}$’s, which corresponds to the larger of the corresponding $\nu$’s. The change in the shape of the resulting spinodal curves illustrates the transition between the two types of phase diagrams shown in Fig. S2, as the spinodal curve starts to protrude through the binodal. [Note that, in regions where $\phi_{a}^{\rm st}(q)<\phi_{a}^{\rm osc}$, the dashed curve $\phi_{a}=\phi_{a}^{\rm osc}$ does not indicate an eigenvalue with vanishing real part because it lies in a region where $\det(W)>0$. However, the solid blue (spinodal) curve does always indicate such an eigenvalue.]
@@ 2 Protrusion of the spinodal through the binodal
Since the binodal curve is defined by the common tangent construction on $\Phi$ of Eqs. (S15,S16), and the function $g_{0}(\rho)=\Phi^{\prime}(R(\rho))$ in (S8), one sees that $g_{0}$ must have two turning points between $\phi_{l}$ and $\phi_{v}$. The equality (S26) that appears in the linear stability analysis is equivalent to $g_{0}^{\prime}=0$. If the spinodal instability is of stationary type, this means that points of inflection of $\Phi$ correspond to spinodal instabilities, as happens in equilibrium. The analog of the critical point in equilibrium occurs when the two turning points coalesce so that $g^{\prime\prime}=g^{\prime}=0$ (stationary point of inflection). At this point, the binodal and spinodal curves are tangent to each other, which is again analogous to equilibrium. It corresponds to a supercritical pitchfork bifurcation. The left panel of Fig. S2 illustrates this case.
@@
D Numerical methods for the hydrodynamic PDEs
@@ 1 Time-dependent solutions
We initiate the scheme with a perturbation around the homogeneous state, $\rho_{\sigma}(x,0)=\phi_{\sigma}+\delta\tilde{\rho}_{\sigma}(x)$ with $\phi_{\pm}=\phi_{a}/2$ and $\phi_{0}=\phi_{p}$. We normalise the perturbation so that $\|(\tilde{\rho}_{+},\tilde{\rho}_{0},\tilde{\rho}_{-})\|_{2}=1$, where $\|u\|_{2}=\left(\int_{0}^{L}|u|^{2}\mathrm{d}x\right)^{1/2}$ is the $L_{2}$ norm. For a random perturbation, we define
@@
We also use the eigenfunctions from linear stability analysis to generate perturbations. In particular, we solve (S22) for $q=2\pi/L$ and select the solution corresponding to the eigenvalue, $\lambda$, with the largest real part and non-negative imaginary part. We define left and right traveling perturbations,
@@ 2 Traveling solutions in the finite system
$(c/L)\varrho_{\sigma}+M_{\sigma}\partial_{z}U_{\sigma}+\sigma\int_{-L/2}^{z}\mathrm{m}(y)\mathrm{d}y=A_{\sigma},$ (S38)
where the mobilities $M_{\sigma}$ and velocities $U_{\sigma}$ are given as in Subsection 1 but replacing $\rho_{\sigma}$ by their traveling frame counterparts $\varrho_{\sigma}$, and $A_{\sigma}$ are integration constants. Additionally, we have the mass constraints
@@
We use the same finite-volume spatial discretization as in Subsection 1, resulting in the discretized equations
$\frac{c}{2L}\left(\varrho_{\sigma,i}+\varrho_{\sigma,i+1}\right)+F_{\sigma,i+1/2}+\sigma\Delta z\sum_{j=0}^{i}\mathrm{m}_{j}$ $=A_{\sigma},$ (S40)
for $i=0,\dots N-1$, where $F_{\sigma,i+1/2}$ is defined in (S33) and mass constraints
@@
The nonlinear system of $3N$ equations (S40)-(S41) is then solved numerically using NonlinearSolve() from the Julia NonlinearSolve.jl package [49], with paramters reltol=1e-8, abstol=1e-8 and maxiters=20. NonlinearSolve() first tries less robust Quasi-Newton methods for more performance and then tries more robust techniques if the faster ones fail. As such, it requires a close enough initial guess for convergence. The first time, we initialize the iterative solver with the long-time ($T=1000$) solution of the time-dependent problem (Subsection 1) with parameters $\phi=0.67,\phi_{a}=0.37$,$\text{Pe}=7.5$, $L=25$,$N=500$. Once we have found a traveling solution to (S40,S41), we use it as an initial condition for the problem with slightly altered parameters: $\phi_{a}^{\prime}=\phi_{a}\pm 0.01$, $\phi_{p}^{\prime}=\phi_{p}\pm 0.01$, as well as larger domains, e.g., $L^{\prime}=2L$ or $N^{\prime}=2N$.
@@ 3 Traveling solutions in large systems
$c\varrho+F&=A_{1},&\qquad F&=-\varrho^{\prime}-\tfrac{\text{Pe}^{2}}{2}(1-\varrho)\partial_{z}[d_{s}(\varrho)\varrho_{a}]\\ c\varrho_{a}+F_{a}&=A_{2},&F_{a}&=-d_{s}(\varrho)\varrho_{a}^{\prime}-\varrho_{a}{\cal D}(\varrho)\varrho^{\prime}-\tfrac{\text{Pe}^{2}}{2}\left[\varrho_{a}s(\varrho)+d_{s}(\varrho)\right]\partial_{z}[d_{s}(\varrho)\varrho_{a}],\end{aligned}$ (S42)
@@
This yields a set of $32N+$ 3 unknowns ($\varrho_{i},\varrho_{a,i},c,A_{1},A_{2}$) and $22N+$ 2 equations [Eqs. (S44) and (S45)]. As in Subsection 2, the additional constraint comes from noting that the problem has translational symmetry; we fix $\varrho_{0}=\phi$.
@@
Figure S3: CP solution to (1). The initial condition is the homogeneous state ($\rho_{\pm}=\phi_{a}/2,\rho_{0}=\phi_{p}$) perturbed by a uniform random perturbation (see Sec. D 1 for details). (a) Density profile at $t=500$. (b) Kymographs showing the spatiotemporal dynamics. Parameters: $\text{Pe}=7.5$, $L=100.0$, $\phi_{a}=0.5$, $\phi_{p}=0.3$, $\Delta x=0.05$.
@@
For the outer1 problem with parameters $\phi$ and $\nu$, we first solve the coexisting phases problem (Section B) given $\nu$. This results in vapor and liquid phases $\phi_{v}(\nu)$ and $\phi_{l}(\nu)$, respectively, as well as individual components $(\varrho_{\sigma})_{v}$ and $(\varrho_{\sigma})_{l}$ in each phase. They satisfy $(\varrho_{\sigma})_{v}\equiv(\varrho_{+},\varrho_{-},\varrho_{0})_{v}=(\tfrac{1}{2}\phi_{a,g},\tfrac{1}{2}\phi_{a,g},\phi_{p,g})$ and similarly for $(\varrho_{\sigma})_{l}$. If $\phi_{v}(\nu)=\phi_{l}(\nu)$, it means there is no inner region and outer1 has no solution (and we revert to outer0). Next, we solve the outer problem (S42) with Dirichlet boundary conditions $\varrho(0)=\phi_{v},\varrho(L)=\phi_{l},\varrho_{a}(0)=\phi_{a,g}$ and $\varrho_{a}(L)=\phi_{a,l}$. In contrast to outer0, we now have $32(N-1)+$ 3 unknowns ($\varrho_{i},\varrho_{a,i}$ for $i=1,\dots,N-1$ and $c,A_{1},A_{2}$) and $12N+$ 1 equations [Eqs. (S44) for $i=0,\dots,N-1$ and one mass constraint (S45)]. The actives volume fraction $\phi_{a}$ is found a posteriori as $\phi_{a}=(1/N)\sum_{i=0}^{N-1}\varrho_{a,i}$.
@@
To obtain a solution to the outer2 problem with two interfaces at a distance $\delta$ and the first one with a tie-line parameter $\nu$, the first step is as in outer1: we solve the coexisting phases problem with $\nu_{1}$ and this gives us Dirichlet boundary conditions at $z=0$ and $z=L$ (as before, if $\phi_{v}(\nu)=\phi_{l}(\nu)$ there is no interface and we revert to outer1). We take the separation between interfaces $\delta$ to be $\delta=(j+1/2)\Delta z$ for some $j$, that is, the second interface lies at the half-grid point $z_{j+1/2}$. To force a second interface between the grid points $z_{j}$ and $z_{j+1}$, the two equations (S44) corresponding to this half-grid point are replaced by three binodal equations (Sec. B)

### v2

@@ Appendix C The method of coexisting phases and the binodal curve
Appendix D Linear stability of homogeneous solutions and the spinodal curve
The homogeneous solution $(\rho,\rho_{a},m)=(\phi,\phi_{a},0)$ is always a solution of Eqs. (S1). We analyze the linear stability of this solution by taking $\rho=\phi+\delta\tilde{\rho}$ , $\rho_{a}=\phi_{a}+\delta\tilde{\rho}_{a}$ , $m=0+\delta\tilde{m}$ with $y$-independent perturbations $\tilde{\rho},\tilde{\rho}_{a},\tilde{m}$. At first order in $\delta$, we obtain
@@ Appendix D Linear stability of homogeneous solutions and the spinodal curve
This matrix is not Hermitian, so its eigenvalues are, in general, complex. However, if the spectrum is complex, then two of the eigenvalues form a complex conjugate pair and the other remains real; this is due to symmetry under $\lambda\mapsto\bar{\lambda}$ and $(A_{1},A_{2},A_{3})\mapsto(\bar{A}_{1},\bar{A}_{2},-\bar{A}_{3})$ in (S46), where $\bar{A}$ denotes the complex conjugate of $A$.
As usual, the homogeneous state is linearly stable if all eigenvalues $\lambda$ have negative real parts. (It is implicit that $\lambda$ depends on $q$; this condition must hold for all $q$.) The resulting instabilities can have several types; we classify them according to the scheme of Ref. [71]. Our system has two conserved densities $\rho,\rho_{a}$ and a nonconserved magnetization $m$. The behavior is controlled by the conserved densities, which restricts the behavior to four of the eight types considered in Ref. [71]. If the dominant eigenvalue of $W$ is real then it is called stationary, else it is called oscillatory; if the instability is initiated by modes with $q\to 0$ then it is called large scale, else it is called small scale. The resulting types are then conserved-Turing (stationary, small scale), Cahn–Hilliard (stationary, large scale), conserved-Hopf (oscillatory, large scale), or conserved-wave (oscillatory, small scale).
Fig. S2 illustrates the range of possible behavior, showing several instabilities that occur on increasing $\phi_{a}$ at fixed total density $\phi$. Row (a) shows the onset of a Cahn-Hilliard instability (stationary, large-scale), row (b) shows the onset of a conserved-Hopf instability (oscillatory, large-scale), and row (c) shows the onset of a conserved-wave instability (oscillatory, small-scale). In the left (right) column, the active volume fraction is sub(super)-critical $\phi_{a}<\phi_{a}^{*}$ where $\max_{q}\text{Re}\lambda<0$ ($\phi_{a}>\phi_{a}^{*}$ where $\max_{q}\text{Re}\lambda>0$). The central column displays the critical active volume fraction $\phi_{a}=\phi_{a}^{*}$ corresponding to $\max_{q}\text{Re}\lambda/q^{2}=0$. We show below that these are the only possible scenarios because the onset of a stationary instability must occur on a large scale.
Figure S2: Eigenvalues of $W$ (S47) at the onset of instability. Active volume fraction $\phi_{a}$ increases from left to right. The critical volume fraction is displayed in the central column. (a) Cahn-Hiliard instability (stationary, large-scale). (b) Conserved-Hopf instability (oscillatory, large-scale). (c) Conserved-wave instability (oscillatory, small-scale).
@@ 1 Spinodal curves
At the boundary of linear stability, then $\text{Re}(\lambda)=0$ for at least one eigenvalue. There are two situations where this can happen
  1. (i)
    $\text{det}(W)=0$, corresponding to a vanishing eigenvalue $\lambda=0$ (stationary instability).
  2. (ii)
    the characteristic polynomial is of the form $(\lambda^{2}+F(W))(\lambda-\text{Tr}(W))=0$, with $F(W)>0$ so that $\lambda=\pm i\sqrt{F(W)}$ is pure imaginary. This situation holds if and only if $F(W)\text{Tr}(W)=\text{det}(W)<0$ (we have always $\text{Tr}(W)<0$ so the final inequality ensures $F(W)>0$).
In the stationary case (i) one may solve $\det W=0$ for $\phi_{a}$ to obtain
@@
In the oscillatory case (ii), one may similarly solve $\text{det}(W)=F(W)\text{Tr}(W)$ to obtain
$\phi_{a}^{{\rm osc}}(q)=\frac{2\phi\left(1+d_{s}(\phi)q^{2}\right)\left[2+2d_{s}(\phi)+d_{s}^{2}(\phi)\text{Pe}^{2}+\left(d_{s}(\phi)+1\right)^{2}q^{2}\right]}{\text{Pe}^{2}\left\{2d_{s}^{3}(\phi)q^{2}+d_{s}^{\prime}(\phi)(\phi-1)\phi(2+q^{2})+d_{s}^{2}(\phi)[2+(\phi-1)q^{2}]+d_{s}(\phi)(\phi-1)[2+(1+d_{s}^{\prime}(\phi)\phi)q^{2}]\right\}}.$ (S51)
@@
[For a finite system, the trial solution (S45) is restricted to $q=2n\pi/L$ with $n\in\mathbb{Z}$ and we should minimize over this discrete set, so $\phi_{a}^{*}$ depends in general on $L$. We consider the limit $L\to\infty$ here, so we take an infimum over $q>0$.] Note also that, if the infimum in (S52) is achieved by $\phi_{a}^{{\rm osc}}(q)$, then it is certain that $\det(W)<0$ at this point, as required for an oscillatory instability: this holds because $\det(W)$ only changes sign at $\phi_{a}=\phi_{a}^{\rm st}(q)>\phi_{a}^{\rm osc}(q)$ and $\det(W)<0$ for $\phi_{a}\to 0$. We also observe from (S50) that $\inf_{q}\phi_{a}^{{\rm st}}(q)=\phi_{a}^{{\rm st}}(0)$, which means that if this instability is of stationary type then it is always large-scale, as already asserted above. On the other hand, the infimum of $\phi_{a}^{{\rm osc}}(q)$ may occur as $q\to 0$ (large-scale oscillatory instability) or at finite $q$ (small-scale oscillatory instability).
@@
For numerical calculations, it is convenient to parameterize the dependence on $\phi_{a}$ in terms of the quantity $\nu=\phi_{p}/(1-\phi)$ defined in Eq. (5). Then $\nu=\frac{\phi}{1-\phi}$ corresponds to a system of purely passive particles and $\nu=0$ to purely active particles. Fig. S3 illustrates the spinodal in the $(\phi,\nu)$-plane. We consider a narrow range of Pe in which both oscillatory and stationary instabilities occur, and we plot separately the curves corresponding to $\phi_{a}=\inf_{q}\phi_{a}^{\rm osc}(q)$ and $\phi_{a}=\inf_{q}\phi_{a}^{\rm st}(q)$. The boundary of linear instability is given by the smaller of these $\phi_{a}$’s, which corresponds to the larger of the corresponding $\nu$’s. The change in the shape of the resulting spinodal curves illustrates the transition between the two types of phase diagrams shown in Fig. S3, as the spinodal curve starts to protrude through the binodal. [Note that, in regions where $\phi_{a}^{\rm st}<\phi_{a}^{\rm osc}$, the dashed curve $\phi_{a}=\phi_{a}^{\rm osc}$ does not indicate an eigenvalue with vanishing real part because it lies in a region where $\det(W)>0$. However, the solid blue (spinodal) curve does always indicate such an eigenvalue.]
@@ 2 Protrusion of the spinodal through the binodal
Since the binodal curve is defined by the common tangent construction on $\Phi$ of Eqs. (S39,S40), and the function $g_{0}(\rho)=\Phi^{\prime}(R(\rho))$ in (S32), one sees that $g_{0}$ must have two turning points between $\phi_{l}$ and $\phi_{v}$. The equality (S50) that appears in the linear stability analysis is equivalent to $g_{0}^{\prime}=0$. If the spinodal instability is of stationary type, this means that points of inflection of $\Phi$ correspond to spinodal instabilities, as happens in equilibrium. The analog of the critical point in equilibrium occurs when the two turning points coalesce so that $g^{\prime\prime}=g^{\prime}=0$ (stationary point of inflection). At this point, the binodal and spinodal curves are tangent to each other, which is again analogous to equilibrium. It corresponds to a supercritical pitchfork bifurcation. The left panel of Fig. S3 illustrates this case.
@@
As a final comment in this section, note that binodal is analytic, and has a quadratic Taylor expansion near its maximum (in the $\nu,\phi$ plane). This maximum is a critical point, and the analytic structure is relevant for the critical behaviour of the model, whose exponents are those of the Gaussian fixed point (of the renormalisation group). This effect can be traced back to the $h$-dependence of the APLG rates, which acts to suppress fluctuations, and enables the exact derivation of the hydrodynamic limit.
3 Small $q$ approximation of the spinodal
Consider (S46) in the $L\gg 1$ limit, corresponding to wavenumber $q\ll 1$. From the first two equations in (S46) we find $\lambda=O(q^{2})$, $A_{1},A_{2}=O(1)$ and $A_{3}=O(q)$. As a result, we can solve for $A_{3}$ in the third equation to obtain
$A_{3}=-\frac{1}{2}iq\text{Pe}\left[d_{s}^{\prime}(\phi)\phi_{a}A_{1}+d_{s}(\phi)A_{2}\right]+O(q^{2}).$ (S53)
The fact that $A_{3}$ (corresponding to the perturbation $\tilde{m}$) drops for $q\ll 1$ is consistent with our asymptotic analysis in Sec. .4, where we find that the magnetization vanishes at leading order. Inserting (S53) into (S46) and rearranging the resulting $2\times 2$ system in terms of $A_{a}\equiv A_{2}$ and $A_{0}=A_{1}-A_{2}$ we find that $\tilde{\lambda}$ for $q\ll 1$ solves
$\widetilde{W}\begin{pmatrix}A_{a}\\ A_{0}\end{pmatrix}=\tilde{\lambda}\begin{pmatrix}A_{a}\\ A_{0}\end{pmatrix},$ (S54)
where $\widetilde{W}=-q^{2}M(H+\frac{\text{Pe}^{2}}{2}G_{\text{NR}})$ and $M,H,G_{\text{NR}}$ coincide with the matrices in (S27,S28) corresponding to the nonreciprocal form of the outer problem. Hence, the matrix of linear stability in the limit $q\ll 1$ coincides, up to the factor $-q^{2}$, with the matrix product $MG$ when writing the outer problem as a cross-diffusion problem (Example 3).
Fig. S4 overlays the spinodal calculated according to (S50) and (S51) with the approximation for small $q$ or outer region solution for $\text{Pe}=7.5$. We find excellent agreement between the approximation and the full calculation, indicating that for this value of Pe, our model always goes unstable through a large-scale instability (either stationary or oscillatory). This is because, for $\text{Pe}\lessapprox 8$, $\inf_{q}\phi_{a}^{{\rm osc}}(q)=\phi_{a}^{{\rm osc}}(0)$.
[…]
Appendix E Numerical methods for the hydrodynamic PDEs
@@ 1 Time-dependent solutions
We initiate the scheme with a perturbation around the homogeneous state, $\rho_{\sigma}(x,0)=\phi_{\sigma}+\delta\tilde{\rho}_{\sigma}(x)$ with $\phi_{\pm}=\phi_{a}/2$ and $\phi_{0}=\phi_{p}$. We normalise the perturbation so that $\|(\tilde{\rho}_{+},\tilde{\rho}_{0},\tilde{\rho}_{-})\|_{2}=1$, where $\|u\|_{2}=\left(\int_{0}^{L}|u|^{2}\mathrm{d}x\right)^{1/2}$ is the $L_{2}$ norm. For a uniformly random perturbation, we define
@@
We also use the eigenfunctions from linear stability analysis to generate perturbations. In particular, we solve (S46) for $q=2\pi/L$ and select the solution corresponding to the eigenvalue, $\lambda$, with the largest real part and nonnegative imaginary part. We define left and right traveling perturbations,
@@ 2 Traveling solutions in the finite system
$-(c/L)\varrho_{\sigma}+M_{\sigma}\partial_{z}U_{\sigma}+\sigma\int_{-L/2}^{z}\mathrm{m}(y)\mathrm{d}y=A_{\sigma},$ (S64)
where $A_{\sigma}$ are integration constants and the mobilities $M_{\sigma}$ and velocities $U_{\sigma}$ are as given in Subsec. 1 but replacing $\rho_{\sigma}$ by their traveling frame counterparts $\varrho_{\sigma}$. Additionally, we have the mass constraints
@@
We use the same finite-volume spatial discretization as in Subsec. 1, resulting in the discretized equations
$-\frac{c}{2L}\left(\varrho_{\sigma,i}+\varrho_{\sigma,i+1}\right)+F_{\sigma,i+1/2}+\sigma\Delta z\sum_{j=0}^{i}\mathrm{m}_{j}$ $=A_{\sigma},$ (S66)
for $i=0,\dots N-1$, where $F_{\sigma,i+1/2}$ is defined in (S59) (but in terms of the travelling frame variables), together with mass constraints
@@
The nonlinear system of $3N$ equations (S66)-(S67) is then solved numerically using NonlinearSolve() from the Julia NonlinearSolve.jl package [63], with parameters reltol=1e-8, abstol=1e-8 and maxiters=20. The solver first tries less robust Quasi-Newton methods for more performance and then tries more robust techniques if the faster ones fail. As such, it requires a close enough initial guess for convergence. The first time, we initialize the iterative solver with the long-time ($T=1000$) solution of the time-dependent problem (Subsec. 1) with parameters $\phi=0.67,\phi_{a}=0.37$,$\text{Pe}=7.5$, $L=25$,$N=500$. Once we have found a traveling solution to (S66,S67), we use it as an initial condition for the problem with slightly altered parameters: $\phi_{a}^{\prime}=\phi_{a}\pm 0.01$, $\phi_{p}^{\prime}=\phi_{p}\pm 0.01$, as well as larger domains, e.g., $L^{\prime}=2L$ or $N^{\prime}=2N$.
@@ 3 Traveling solutions in large systems
$-(c/L)\varrho+F&=A_{1},&\qquad F&=-\varrho^{\prime}-\tfrac{\text{Pe}^{2}}{2}(1-\varrho)\partial_{z}[d_{s}(\varrho)\varrho_{a}]\\ -(c/L)\varrho_{a}+F_{a}&=A_{2},&F_{a}&=-d_{s}(\varrho)\varrho_{a}^{\prime}-\varrho_{a}{\cal D}(\varrho)\varrho^{\prime}-\tfrac{\text{Pe}^{2}}{2}\left[\varrho_{a}s(\varrho)+d_{s}(\varrho)\right]\partial_{z}[d_{s}(\varrho)\varrho_{a}],\end{aligned}$ (S68)
@@
This yields a set of $32N+$ 3 unknowns ($\varrho_{i},\varrho_{a,i},c,A_{1},A_{2}$) and $22N+$ 2 equations [Eqs. (S70) and (S71)]. As in Subsec. 2, the additional constraint comes from noting that the problem has translational symmetry; we fix $\varrho_{0}=\phi$.
@@
Figure S5: CP solution to (1). The initial condition is the homogeneous state ($\rho_{\pm}=\phi_{a}/2,\rho_{0}=\phi_{p}$) perturbed by a uniform random perturbation (see Sec. 1 for details). (a) Density profile at $t=4000$. (b) Kymographs showing the spatiotemporal dynamics. Parameters: $\text{Pe}=7.5$, $L=100.0$, $\phi_{a}=0.5$, $\phi_{p}=0.3$, $\Delta x=0.05$.
Figure S6: Two solutions, both obtained from Eq. (8) with the same parameters. (a) T solution. (b) PS solution. The top panels display the solution profile as a function of $z/L$. The bottom panels display the profile (solid black) overlayed on phase diagrams spanned by $\phi_{a}$ and $\phi_{p}$. The Spinodal (Blue) encloses the region of linear stability of homogeneous solutions, and the Binodal (Red) encloses the region of Phase separation. Black crosses mark the volume fraction. Parameters: $\text{Pe}=7.5$, $\phi_{a}=0.4608$, $\phi_{p}=0.2201$, $N=1024$
Figure S7: (a) Sketch of T solution with positive velocity, showing density of active and passive particles as a function of position (see Fig. 4 of main text for a similar solution with negative velocity). The red arrows indicate the self-propulsive force acting on the active particles. The blue arrows indicate forces applied by the active particles on the passive ones (via the cross-diffusion terms in the equation of motion). This generates the positive propagation. (b) Sketch of a CP solution at three different times in a large system (see Fig. 3 of main text for a similar solution in a smaller system). Initially the CP solution is well-approximated by a combination of two T solutions with opposite velocities of equal magnitude. The two propagating clusters approach each other and eventually collide. After the collision process one again sees a combination of two T solutions which travel away from each other. The cycle will eventually repeat due to periodic boundaries.
@@
For the outer1 problem with parameters $\phi$ and $\nu$, we first solve the coexisting phases problem (Sec. C) given $\nu$. This results in vapor and liquid phases $\phi_{v}(\nu)$ and $\phi_{l}(\nu)$, respectively, as well as individual components $(\varrho_{\sigma})_{v}$ and $(\varrho_{\sigma})_{l}$ in each phase. They satisfy $(\varrho_{\sigma})_{v}\equiv(\varrho_{+},\varrho_{-},\varrho_{0})_{v}=(\tfrac{1}{2}\phi_{a,g},\tfrac{1}{2}\phi_{a,g},\phi_{p,g})$ and similarly for $(\varrho_{\sigma})_{l}$. If $\phi_{v}(\nu)=\phi_{l}(\nu)$, it means there is no inner region and outer1 has no solution (and we revert to outer0). Next, we solve the outer problem (S68) with Dirichlet boundary conditions $\varrho(0)=\phi_{v},\varrho(L)=\phi_{l},\varrho_{a}(0)=\phi_{a,g}$ and $\varrho_{a}(L)=\phi_{a,l}$. In contrast to outer0, we now have $32(N-1)+$ 3 unknowns ($\varrho_{i},\varrho_{a,i}$ for $i=1,\dots,N-1$ and $c,A_{1},A_{2}$) and $12N+$ 1 equations [Eqs. (S70) for $i=0,\dots,N-1$ and one mass constraint (S71)]. The actives volume fraction $\phi_{a}$ is found a posteriori as $\phi_{a}=(1/N)\sum_{i=0}^{N-1}\varrho_{a,i}$.
@@
To obtain a solution to the outer2 problem with two interfaces at a distance $\delta$ and the first one with a tie-line parameter $\nu$, the first step is as in outer1: we solve the coexisting phases problem with $\nu$ and this gives us Dirichlet boundary conditions at $z=0$ and $z=L$ (as before, if $\phi_{v}(\nu)=\phi_{l}(\nu)$ there is no interface and we revert to outer1). We take the separation between interfaces $\delta$ to be $\delta=(j+1/2)\Delta z$ for some $j$, that is, the second interface lies at the half-grid point $z_{j+1/2}$. To force a second interface between the grid points $z_{j}$ and $z_{j+1}$, the two equations (S70) corresponding to this half-grid point are replaced by three binodal equations (Sec. C)